# The Identity Problem in the special affine group of $\mathbb{Z}^{2}$

### v5

@@ 1. Introduction
Apart from their obvious use in determining structural properties of a semigroup [^1] For example, given a decision procedure for the Group Problem, one can compute a generating set for the group of units (set of of invertible elements) of a finitely generated semigroup $\langle\mathcal{G}\rangle$ [Don24]., the Identity Problem and the Group Problem are closely related to [^2] In fact, decidability of Semigroup Membership subsumes decidability of the Identity Problem and the Group Problem. the more difficult Semigroup Membership problem. Usually, the solution to the Identity Problem is the most essential special case on the way to building an algorithm for the Semigroup Membership problem. We also point out that there are significantly more available algorithms for groups then there are for semigroups, therefore performing preliminary checks using the Group Problem can help decide Semigroup Membership in many special cases (see Section 6).
All four algorithmic problems remain undecidable even for matrices of small dimensions. The Special Linear group of dimension $n$, denoted by $\mathsf{SL}(n,\mathbb{Z})$, is defined as the group of $n\times n$ integer matrices with determinant one. Mikhailova famously showed undecidability of the Group Membership problem (and hence also Semigroup Membership) in $\mathsf{SL}(4,\mathbb{Z})$ [Mik66]. Later, Bell and Potapov showed undecidability of the Identity Problem and the Group Problem in $\mathsf{SL}(4,\mathbb{Z})$ [BP10]. Both undecidability results stem from the fact that $\mathsf{SL}(4,\mathbb{Z})$ contains as a subgroup a direct product of two non-abelian free groups. In dimension two, the Semigroup Membership problem in $\mathsf{SL}(2,\mathbb{Z})$ was shown to be decidable in EXPSPACE by Choffrut and Karhumäki [CK05]; Group Membership was shown to be in PTIME by Lohrey [Loh23]; the Identity Problem and the Group Problem were shown to be NP-complete by Bell, Hirvensalo and Potapov [BHP17]. It remains an intricate open problem whether any of these four algorithmic problems is decidable in $\mathsf{SL}(3,\mathbb{Z})$. Nevertheless, Ko, Niskanen and Potapov [KNP18a] recently showed that $\mathsf{SL}(3,\mathbb{Z})$ cannot embed pairs of words over an alphabet of size two, suggesting that all four problems in $\mathsf{SL}(3,\mathbb{Z})$ might be decidable. The following table 1 summarizes the state of art as well as our result, the group $\mathsf{SA}(2,\mathbb{Z})$ will be introduced in the next subsection.
@@ 2. Preliminaries
For an arbitrary matrix $A\in\mathsf{SL}(2,\mathbb{Z})$, an invariant subspace of $A$ is a $\mathbb{C}$-linear subspace $V$ of $\mathbb{C}^{2}$ such that $AV=V$. If the eigenvalues of $A$ are reals, then one can suppose that the invariant spaces of $A$ are subspaces of $\mathbb{R}^{2}$. Denote by $\operatorname{Lat}(A)$ the set of dimension one invariant subspaces of $A$. It is easy to see that if $A\notin\{I,-I\}$, then $\operatorname{Lat}(A)$ has one or two elements.
@@
Given an arbitrary group $G$ with neutral element $I$. An element $T\in G$ is called torsion if $T\neq I$ and $T^{m}=I$ for some $m>1$. A group is called torsion-free if it does not contain a torsion element. In particular, a free group is torsion-free.
@@
Let $A=\begin{pmatrix}a&b\\ c&d\\ \end{pmatrix}$ be a matrix in $\mathsf{SL}(2,\mathbb{Z})$. The characteristic polynomial of $A$ is $f(X)=X^{2}-(a+d)X+(ad-bc)=X^{2}-(a+d)X+1$. Consider the five following cases.
@@
    In this case, either $A=-I$, or $A\overset{\mathbb{Q}}{\sim}\begin{pmatrix}-1&1\\ 0&-1\\ \end{pmatrix}$. In the second situation, we call $A$ a twisted inversion. In particular, if $A$ is a twisted inversion, then $(A+I)^{2}=0$, and $A^{2}$ is a shear, and $\operatorname{Lat}(A)$ has exactly one element.
@@
    In this case, $A\overset{\mathbb{R}}{\sim}\begin{pmatrix}\lambda&0\\ 0&\lambda^{-1}\\ \end{pmatrix}$, where $\lambda$ is the root of $f(X)$ such that $|\lambda|\geq 1$. Furthermore, $\lambda$ is real. In this case, we call $A$ a scale. If $\lambda>0$, we call $A$ a positive scale; if $\lambda<0$, we call $A$ an inverting scale. In both cases, $\operatorname{Lat}(A)$ has two elements, one element is the invariant space corresponding to the eigenvalue $\lambda$, and is called the stretching direction; the other element is the invariant space corresponding to the eigenvalue $\lambda^{-1}$, and is called the compressing direction. See Figure 2 for an illustration.
@@ 3. Overview of decision procedures
We now first focus on solving the Group Problem, because by Lemma 2, decidability of the Group Problem will imply decidability of the Identity Problem. Fix a set $\mathcal{G}\coloneqq\{(A_{1},\boldsymbol{a}_{1}),\ldots,(A_{K},\boldsymbol{a}_{K})\}$ of elements in $\mathsf{SA}(2,\mathbb{Z})$. The following lemma shows that, if the semigroup $G\coloneqq\langle A_{1},\ldots,A_{K}\rangle$ is not a group, then $\langle\mathcal{G}\rangle$ is also not a group. Therefore, to decide whether $\langle\mathcal{G}\rangle$ is a group, we can focus on the case where $G$ is a group.
@@
In case of $G$ being virtually solvable, we will prove the following Proposition 7 which refines the Tits alternative. In particular, it shows that virtually solvable subgroups of $\mathsf{SL}(2,\mathbb{Z})$ have relatively simple structure: it is either trivial, or it contains a torsion element, or it is infinite cyclic.
@@
Next we solve the Identity Problem. Note that by Lemma 2, $I\in\langle\mathcal{G}\rangle$ if and only if there exists a non-empty subset $\mathcal{H}$ of $\mathcal{G}$ such that the semigroup $\langle\mathcal{H}\rangle$ is a group. Therefore, it suffices to guess a non-empty subset $\mathcal{H}$ of $\mathcal{G}$, and check whether $\langle\mathcal{H}\rangle$ is a group. This can be done in NP using the above procedure of the Group Problem.
@@
    If $G$ is virtually free, decide which case of Proposition 7 is true using Lemma 14.
@@
      If $G$ is trivial, decide whether $n_{1}\boldsymbol{a}_{1}+\cdots n_{K}\boldsymbol{a}_{K}=\boldsymbol{0}$ has a solution over $\mathbb{Z}_{>0}^{K}$. If yes, return True, otherwise return False. (Proposition 15)
@@ 4. Non-abelian free subgroup
In other words, the acute angle $\theta$ between $\boldsymbol{y}$ and $V$ satisfies $1-\cos\theta<\varepsilon$. Also, $\boldsymbol{y}$ and $\boldsymbol{x}$ lie in same cone out of the four cut out by $V$ and $W$. See Figure 4 for an illustration.
@@
    If $V$ is a stretching direction. In this case, $|\lambda|>1$. Let $m>0$ be a positive integer, and let $w\coloneqq(A,\boldsymbol{a})^{m-1}(B,\boldsymbol{b})^{m-1}$. Consider the product
@@
    If $V$ is a compressing direction. In this case, $|\lambda|<1$. Let $m>0$ be a positive integer, and let $w\coloneqq(B,\boldsymbol{b})^{m}(A,\boldsymbol{a})^{m}$. Consider the product
@@
    If $\mu c>0$. Let $m>0$ be a positive integer, and let $w\coloneqq(A,\boldsymbol{a})^{m-1}(B,\boldsymbol{b})^{m-1}$. Consider the product
@@
    When $m\rightarrow\infty$, this expression tends towards $0+1=1$, because $\mu c>0$ and $\boldsymbol{v}\neq\boldsymbol{0}$. Hence, for a large enough odd integer $m$, we have $1-\frac{\boldsymbol{v}^{\top}\boldsymbol{y}}{\|\boldsymbol{v}\|\|\boldsymbol{y}\|}<\varepsilon.$
@@
    If $\mu c<0$. Let $m>0$ be a positive integer, and let $w\coloneqq(B,\boldsymbol{b})^{m}(A,\boldsymbol{a})^{m}$. Consider the product
@@
    When $m\rightarrow\infty$, this expression tends towards $0+1=1$, because $\mu c<0$ and $\boldsymbol{v}\neq\boldsymbol{0}$. Hence, for a large enough odd integer $m$, we have $1-\frac{\boldsymbol{v}^{\top}\boldsymbol{y}}{\|\boldsymbol{v}\|\|\boldsymbol{y}\|}<\varepsilon.$
@@
Case 1. Both $A$ and $B$ are positive scales. Denote $Y=A^{-1}B^{-1}\in\langle A_{1},\ldots,A_{K}\rangle$, we have $AYB=I$. Let $\boldsymbol{y}\in\mathbb{Z}^{2}$ be such that $(Y,\boldsymbol{y})\in\langle\mathcal{G}\rangle$, and let $\boldsymbol{x}\in\mathbb{Z}^{2}$ be such that $(A,\boldsymbol{a})(Y,\boldsymbol{y})(B,\boldsymbol{b})=(I,\boldsymbol{x})$. If $\boldsymbol{x}=\boldsymbol{0}$, then $(I,\boldsymbol{0})$ can be represented as a full-image word over $\mathcal{G}$, since $(A,\boldsymbol{a})$ can. In this case, $\langle\mathcal{G}\rangle$ is a group by Lemma 1. Otherwise, there are two subcases. Starting from any one-dimension subspace $S$, rotate $S$ counter-clockwise and consider its sequence of encounters with $\operatorname{Lat}(A)$ and $\operatorname{Lat}(B)$ before finishing a full cycle. Either the sequence of encounters is a cyclic permutation of $(\operatorname{Lat}(A),\operatorname{Lat}(A),\operatorname{Lat}(B),\operatorname{Lat}(B))$, or it is a cyclic permutation of $(\operatorname{Lat}(A),\operatorname{Lat}(B),\operatorname{Lat}(A),\operatorname{Lat}(B))$.
@@
    First consider the case of $(\operatorname{Lat}(A)$, $\operatorname{Lat}(A)$, $\operatorname{Lat}(B)$, $\operatorname{Lat}(B))$. An illustration for the proof of this case is shown in Figure 8. Denote by $V_{A},W_{A},V_{B},W_{B}$ these subspaces in this order. Since $\boldsymbol{x}\neq\boldsymbol{0}$, either $\boldsymbol{x}\notin V_{A}\cup W_{A}$, or $\boldsymbol{x}\notin V_{B}\cup W_{B}$. Without loss of generality suppose $\boldsymbol{x}\notin V_{B}\cup W_{B}$.
@@
    Let $\varepsilon>0$. Apply Lemma 9 to $\varepsilon$, to the elements $(A^{\prime},\boldsymbol{a}^{\prime})\coloneqq(A,\boldsymbol{a})(Y,\boldsymbol{y})$ and $(B,\boldsymbol{b})$, and to the subspaces $V_{B},W_{B}$ of $\operatorname{Lat}(B)=\operatorname{Lat}(A^{\prime})$. (Note that $A^{\prime}=B^{-1}$, so $\operatorname{Lat}(A^{\prime})=\operatorname{Lat}(B)$.) Since $(A^{\prime},\boldsymbol{a}^{\prime})(B,\boldsymbol{b})=(A,\boldsymbol{a})(Y,\boldsymbol{y})(B,\boldsymbol{b})=(I,\boldsymbol{x})$, Lemma 9 shows there exists an element $w\in\langle\mathcal{G}\rangle$ such that $(A^{\prime},\boldsymbol{a}^{\prime})\cdot w\cdot(B,\boldsymbol{b})=(I,\boldsymbol{x}_{1})$, with
@@
    Here, $\boldsymbol{v}_{B}$ is a non-zero vector in $V_{B}$. In other words, when $\varepsilon$ is sufficiently small, the acute angle between $\boldsymbol{x}_{1}$ and $V_{B}$ is also sufficiently small. Also, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}$ lie in same cone out of the four cut by $V_{B}$ and $W_{B}$. This gives us an element $(Y_{1},\boldsymbol{y}_{1})\coloneqq(Y,\boldsymbol{y})\cdot w\in\langle\mathcal{G}\rangle$ such that $(A,\boldsymbol{a})(Y_{1},\boldsymbol{y}_{1})(B,\boldsymbol{b})=(I,\boldsymbol{x}_{1})$ with $\boldsymbol{x}_{1}$ satisfying (4), see Figure 8.
@@
    $01-\frac{|\boldsymbol{v}_{A}^{\top}\boldsymbol{x}_{2}|}{\|\boldsymbol{v}_{A}\|\|\boldsymbol{x}_{2}\|}<\varepsilon,\;\left(\boldsymbol{x}_{1}\right)_{V_{A}}^{\top}\boldsymbol{x}_{V_{A}}>0,\;\left(\boldsymbol{x}_{1}\right)_{W_{A}}^{\top}\cdot\boldsymbol{x}_{W_{A}}>$ 0 . (5)
    Here, $\boldsymbol{v}_{A}$ is a non-zero vector in $V_{A}$. In other words, when $\varepsilon$ is sufficiently small, the acute angle between $\boldsymbol{x}_{2}$ and $V_{A}$ is also sufficiently small. Hence, one can take $\varepsilon$ such that $\boldsymbol{x}_{2}$ and $-\boldsymbol{x}_{1}$ lie in the same cone out of the four cut by $V_{B}$ and $W_{B}$. Furthermore, $\boldsymbol{x}_{2}$ and $\boldsymbol{x}_{1}$ lie in same cone out of the four cut by $V_{A}$ and $W_{A}$.
@@
    Next consider the case $(\operatorname{Lat}(A)$, $\operatorname{Lat}(B)$, $\operatorname{Lat}(A)$, $\operatorname{Lat}(B))$. Denote by $V_{A},V_{B},W_{A},W_{B}$ these subspaces in this order. Without loss of generality suppose $\boldsymbol{x}\notin V_{B}\cup W_{B}$. The strategy is exactly the same as the previous case, see Figure 8. However, in the present case, we need to apply Lemma 9 for a total of six times, where $V$ will be the subspaces $V_{B},W_{A},W_{B},V_{A},V_{B},W_{A}$ respectively.
@@
    Similarly, since the element $(A,\boldsymbol{a})$ (and hence $(I,\boldsymbol{0})$) can be represented as a full-image word over $\mathcal{G}$, the semigroup $\langle\mathcal{G}\rangle$ is a group by Lemma 1.
@@
Without loss of generality, let $A$ be the shear and $B$ be the positive scale. Starting from any one-dimension subspace $S$, rotate $S$ counter-clockwise and consider its sequence of encounters with $\operatorname{Lat}(A)$ and $\operatorname{Lat}(B)$ before finishing a full cycle. The sequence of encounters must a cyclic permutation of $(\operatorname{Lat}(A),\operatorname{Lat}(B),\operatorname{Lat}(B))$. Denote by $V_{A},V_{B},W_{B}$ be these subspaces in this order. Without loss of generality suppose $\boldsymbol{x}\notin V_{B}\cup W_{B}$, the case where $\boldsymbol{x}\notin V_{A}$ is analogous.
For a small enough $\varepsilon$, apply Lemma 9 to $\boldsymbol{x}$ to obtain $\boldsymbol{x}_{1}$ that is sufficiently close to $V_{B}$; then apply Lemma 10 to $\boldsymbol{x}_{1}$ to obtain $\boldsymbol{x}_{2}$ that is sufficiently close to $V_{A}$ and lies in the same cone cut by $V_{B}$ and $W_{B}$ as $-\boldsymbol{x}_{1}$; then apply Lemma 9 to $\boldsymbol{x}_{2}$ to obtain $\boldsymbol{x}_{3}$ that is sufficiently close to $V_{B}$ and lies in a different halfspace cut by $V_{A}$ than $\boldsymbol{x}_{2}$; finally, apply Lemma 10 to $\boldsymbol{x}_{3}$ to obtain $\boldsymbol{x}_{4}$ that is sufficiently close to $V_{A}$ and lies in the same cone cut by $V_{B}$ and $W_{B}$ as $-\boldsymbol{x}_{3}$. In this way, we obtain $(I,\boldsymbol{x}_{1}),(I,\boldsymbol{x}_{2}),(I,\boldsymbol{x}_{3}),(I,\boldsymbol{x}_{4})$ with $\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{4}$ generating $\mathbb{Q}^{2}$ as a $\mathbb{Q}_{\geq 0}$-cone. Hence, there exist positive integers $n_{1},\ldots,n_{4}$ such that

### v6

@@ 1. Introduction
Apart from their obvious use in determining structural properties of a semigroup [^1] For example, given a decision procedure for the Group Problem, one can compute a generating set for the group of units (set of of invertible elements) of a finitely generated semigroup $\langle\mathcal{G}\rangle$ [Don24]., the Identity Problem and the Group Problem are closely related to [^2] In fact, decidability of Semigroup Membership subsumes decidability of the Identity Problem and the Group Problem. the more difficult Semigroup Membership problem. Usually, the solution to the Identity Problem is the most essential special case on the way to building an algorithm for the Semigroup Membership problem. We also point out that there are significantly more available algorithms for groups than there are for semigroups, therefore performing preliminary checks using the Group Problem can help decide Semigroup Membership in many special cases (see Section 6).
All four algorithmic problems remain undecidable even for matrices of small dimensions. The Special Linear group of dimension $n$, denoted by $\mathsf{SL}(n,\mathbb{Z})$, is defined as the group of $n\times n$ integer matrices with determinant one. Mikhailova famously showed undecidability of the Group Membership problem (and hence also Semigroup Membership) in $\mathsf{SL}(4,\mathbb{Z})$ [Mik66]. Later, Bell and Potapov showed undecidability of the Identity Problem and the Group Problem in $\mathsf{SL}(4,\mathbb{Z})$ [BP10]. Both undecidability results stem from the fact that $\mathsf{SL}(4,\mathbb{Z})$ contains as a subgroup a direct product of two non-abelian free groups. In dimension two, the Semigroup Membership problem in $\mathsf{SL}(2,\mathbb{Z})$ was shown to be decidable in EXPSPACE by Choffrut and Karhumäki [CK05]; Group Membership was shown to be in PTIME by Lohrey [Loh23]; the Identity Problem and the Group Problem were shown to be NP-complete by Bell, Hirvensalo and Potapov [BHP17]. Here, elements of $\mathsf{SL}(2,\mathbb{Z})$ are represented using matrices with binary encoded entries. It remains an intricate open problem whether any of these four algorithmic problems is decidable in $\mathsf{SL}(3,\mathbb{Z})$. Nevertheless, Ko, Niskanen and Potapov [KNP18a] recently showed that $\mathsf{SL}(3,\mathbb{Z})$ cannot embed pairs of words over an alphabet of size two, suggesting that all four problems in $\mathsf{SL}(3,\mathbb{Z})$ might be decidable. The following table 1 summarizes the state of art as well as our result, the group $\mathsf{SA}(2,\mathbb{Z})$ will be introduced in the next subsection.
@@ 2. Preliminaries
For an arbitrary matrix $A\in\mathsf{SL}(2,\mathbb{Z})$, an invariant subspace of $A$ is a $\mathbb{C}$-linear subspace $V$ of $\mathbb{C}^{2}$ such that $AV=V$. If the eigenvalues of $A$ are reals, then one can suppose that the invariant spaces of $A$ are subspaces of $\mathbb{R}^{2}$. Denote by $\operatorname{Lat}(A)$ the set of dimension one invariant subspaces of $A$. It is easy to see that if $A\notin\{I,-I\}$, then all the eigenspaces of $A$ have dimension one, so $\operatorname{Lat}(A)$ has one or two elements.
@@
Let $G$ be an arbitrary group with neutral element $I$. An element $T\in G$ is called torsion if $T\neq I$ and $T^{m}=I$ for some $m>1$. A group is called torsion-free if it does not contain a torsion element. In particular, a free group is torsion-free.
@@
Let $A=\begin{pmatrix}a&b\\ c&d\\ \end{pmatrix}$ be a matrix in $\mathsf{SL}(2,\mathbb{Z})$. The characteristic polynomial of $A$ is $f(X)=X^{2}-(a+d)X+(ad-bc)=X^{2}-(a+d)X+1$. Consider the Jordan Normal form of the matrix $A$ in the five following cases.
@@
    In this case, either $A=-I$, or $A\overset{\mathbb{Q}}{\sim}\begin{pmatrix}-1&1\\ 0&-1\\ \end{pmatrix}$. In the second situation, we call $A$ a twisted inversion. In particular, if $A$ is a twisted inversion, then $(A+I)^{2}=0$, $A^{2}$ is a shear and $\operatorname{Lat}(A)$ has exactly one element.
@@
    In this case, $A\overset{\mathbb{R}}{\sim}\begin{pmatrix}\lambda&0\\ 0&\lambda^{-1}\\ \end{pmatrix}$, where $\lambda$ is the root of $f(X)$ such that $|\lambda|>1$. Furthermore, $\lambda$ is real. In this case, we call $A$ a scale. If $\lambda>0$, we call $A$ a positive scale; if $\lambda<0$, we call $A$ an inverting scale. In both cases, $\operatorname{Lat}(A)$ has two elements, one element is the invariant space corresponding to the eigenvalue $\lambda$, and is called the stretching direction; the other element is the invariant space corresponding to the eigenvalue $\lambda^{-1}$, and is called the compressing direction. See Figure 2 for an illustration.
Figure 1. Illustration for a shear.
@@ 3. Overview of decision procedures
We now first focus on solving the Group Problem, because by Lemma 2, decidability of the Group Problem will imply decidability of the Identity Problem. Fix a set $\mathcal{G}\coloneqq\{(A_{1},\boldsymbol{a}_{1}),\ldots,(A_{K},\boldsymbol{a}_{K})\}$ of elements in $\mathsf{SA}(2,\mathbb{Z})$. The following lemma shows that we only need to consider the case where $G\coloneqq\langle A_{1},\ldots,A_{K}\rangle$ is a group.
@@
In case of $G$ being virtually solvable, we will prove the following Proposition 7 which refines the Tits alternative. [^5] The semigroup $\langle\mathcal{G}\rangle$ is contained in the semidirect product $\mathbb{Z}^{2}\rtimes G\leq\mathsf{SA}(2,\mathbb{Z})$, which is virtually solvable if $G$ is virtually solvable. A recent result by Bodart and Dong [BD24] shows that the Identity Problem and the Group Problem are decidable in all virtually solvable matrix groups over algebraic numbers. This notably includes virtually solvable subgroups of $\mathsf{SA}(2,\mathbb{Z})$. However, it is not clear whether the algorithm of Bodart and Dong runs in elementary time, let alone PTIME. In particular, it shows that a virtually solvable subgroup of $\mathsf{SL}(2,\mathbb{Z})$ has a relatively simple structure: it is either trivial, or it contains a torsion element, or it is infinite cyclic.
@@
Next we solve the Identity Problem. Note that by Lemma 2, $I\in\langle\mathcal{G}\rangle$ if and only if there exists a non-empty subset $\mathcal{H}$ of $\mathcal{G}$ such that the semigroup $\langle\mathcal{H}\rangle$ is a group. Therefore, it suffices to guess a non-empty subset $\mathcal{H}$ of $\mathcal{G}$, and check whether $\langle\mathcal{H}\rangle$ is a group. This can be done in NP using the above procedure for the Group Problem.
@@
    If $G$ is virtually solvable, decide which case of Proposition 7 is true using Lemma 14.
@@
      If $G$ is trivial, decide whether $n_{1}\boldsymbol{a}_{1}+\cdots+n_{K}\boldsymbol{a}_{K}=\boldsymbol{0}$ has a solution over $\mathbb{Z}_{>0}^{K}$. If yes, return True, otherwise return False. (Proposition 15)
@@ 4. Non-abelian free subgroup
In other words, the acute angle $\theta$ between $\boldsymbol{y}$ and $V$ satisfies $1-\cos\theta<\varepsilon$. Also, $\boldsymbol{y}$ and $\boldsymbol{x}$ lie in the same cone out of the four cut out by $V$ and $W$. See Figure 4 for an illustration.
@@
Since $(A,\boldsymbol{a})\cdot(B,\boldsymbol{b})=(I,\boldsymbol{x})$, we have $B=A^{-1}$ and $\boldsymbol{x}=A\boldsymbol{b}+\boldsymbol{a}$. In particular, we have $\operatorname{Lat}(A)=\operatorname{Lat}(B)$.
@@
    $V$ is a stretching direction. In this case, $|\lambda|>1$. Let $m>0$ be a positive integer, and let $w\coloneqq(A,\boldsymbol{a})^{m-1}(B,\boldsymbol{b})^{m-1}$. Consider the product
@@
    $V$ is a compressing direction. In this case, $|\lambda|<1$. Let $m>0$ be a positive integer, and let $w\coloneqq(B,\boldsymbol{b})^{m}(A,\boldsymbol{a})^{m}$. Consider the product
@@
    $\mu c>0$. Let $m>0$ be a positive integer, and let $w\coloneqq(A,\boldsymbol{a})^{m-1}(B,\boldsymbol{b})^{m-1}$. Consider the product
@@
    When $m\rightarrow\infty$, this expression tends towards $0+1=1$, because $\mu c>0$ and $\boldsymbol{v}\neq\boldsymbol{0}$. Hence, for a large enough integer $m$, we have $1-\frac{\boldsymbol{v}^{\top}\boldsymbol{y}}{\|\boldsymbol{v}\|\|\boldsymbol{y}\|}<\varepsilon.$
@@
    $\mu c<0$. Let $m>0$ be a positive integer, and let $w\coloneqq(B,\boldsymbol{b})^{m}(A,\boldsymbol{a})^{m}$. Consider the product
@@
    When $m\rightarrow\infty$, this expression tends towards $0+1=1$, because $\mu c<0$ and $\boldsymbol{v}\neq\boldsymbol{0}$. Hence, for a large enough integer $m$, we have $1-\frac{\boldsymbol{v}^{\top}\boldsymbol{y}}{\|\boldsymbol{v}\|\|\boldsymbol{y}\|}<\varepsilon.$
@@
Case 1. Both $A$ and $B$ are positive scales. Denote $Y=A^{-1}B^{-1}\in\langle A_{1},\ldots,A_{K}\rangle$, we have $AYB=I$. Let $\boldsymbol{y}\in\mathbb{Z}^{2}$ be such that $(Y,\boldsymbol{y})\in\langle\mathcal{G}\rangle$, and let $\boldsymbol{x}\in\mathbb{Z}^{2}$ be such that $(A,\boldsymbol{a})(Y,\boldsymbol{y})(B,\boldsymbol{b})=(I,\boldsymbol{x})$. If $\boldsymbol{x}=\boldsymbol{0}$, then $(I,\boldsymbol{0})$ can be represented as a full-image word over $\mathcal{G}$, since $(A,\boldsymbol{a})$ can. In this case, $\langle\mathcal{G}\rangle$ is a group by Lemma 1. Otherwise, there are two subcases. Starting from any one-dimensional subspace $S$, rotate $S$ counter-clockwise and consider its sequence of encounters with $\operatorname{Lat}(A)$ and $\operatorname{Lat}(B)$ before finishing a half circle. Either the sequence of encounters is a cyclic permutation of $(\operatorname{Lat}(A),\operatorname{Lat}(A),\operatorname{Lat}(B),\operatorname{Lat}(B))$, or it is a cyclic permutation of $(\operatorname{Lat}(A),\operatorname{Lat}(B),\operatorname{Lat}(A),\operatorname{Lat}(B))$.
@@
    First consider the case of $(\operatorname{Lat}(A)$, $\operatorname{Lat}(A)$, $\operatorname{Lat}(B)$, $\operatorname{Lat}(B))$. An illustration for the proof of this case is shown in Figure 8. Denote by $V_{A},W_{A},V_{B},W_{B}$ these subspaces in this order. Since $\boldsymbol{x}\neq\boldsymbol{0}$, we have $\boldsymbol{x}\notin V_{A}\cup W_{A}$ or $\boldsymbol{x}\notin V_{B}\cup W_{B}$. Without loss of generality suppose $\boldsymbol{x}\notin V_{B}\cup W_{B}$.
@@
    Let $\varepsilon>0$. Apply Lemma 9 to $\varepsilon$, to the elements $(A^{\prime},\boldsymbol{a}^{\prime})\coloneqq(A,\boldsymbol{a})(Y,\boldsymbol{y})$ and $(B,\boldsymbol{b})$, and to the subspaces $V_{B},W_{B}$ of $\operatorname{Lat}(B)=\operatorname{Lat}(A^{\prime})$. (Note that $A^{\prime}=B^{-1}$, so $A^{\prime}$ is a positive scale and $\operatorname{Lat}(A^{\prime})=\operatorname{Lat}(B)$.) Since $(A^{\prime},\boldsymbol{a}^{\prime})(B,\boldsymbol{b})=(A,\boldsymbol{a})(Y,\boldsymbol{y})(B,\boldsymbol{b})=(I,\boldsymbol{x})$, Lemma 9 shows there exists an element $w\in\langle\mathcal{G}\rangle$ such that $(A^{\prime},\boldsymbol{a}^{\prime})\cdot w\cdot(B,\boldsymbol{b})=(I,\boldsymbol{x}_{1})$, with
@@
    Here, $\boldsymbol{v}_{B}$ is a non-zero vector in $V_{B}$. In other words, when $\varepsilon$ is sufficiently small, the acute angle between $\boldsymbol{x}_{1}$ and $V_{B}$ is also sufficiently small. Also, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}$ lie in the same cone out of the four cut by $V_{B}$ and $W_{B}$. This gives us an element $(Y_{1},\boldsymbol{y}_{1})\coloneqq(Y,\boldsymbol{y})\cdot w\in\langle\mathcal{G}\rangle$ such that $(A,\boldsymbol{a})(Y_{1},\boldsymbol{y}_{1})(B,\boldsymbol{b})=(I,\boldsymbol{x}_{1})$ with $\boldsymbol{x}_{1}$ satisfying (4), see Figure 8.
@@
    $01-\frac{|\boldsymbol{v}_{A}^{\top}\boldsymbol{x}_{2}|}{\|\boldsymbol{v}_{A}\|\|\boldsymbol{x}_{2}\|}<\varepsilon,\;\left(\boldsymbol{x}_{2}\right)_{V_{A}}^{\top}\boldsymbol{x}_{V_{A}}>0,\;\left(\boldsymbol{x}_{2}\right)_{W_{A}}^{\top}\cdot\boldsymbol{x}_{W_{A}}>$ 0 . (5)
    Here, $\boldsymbol{v}_{A}$ is a non-zero vector in $V_{A}$. In other words, when $\varepsilon$ is sufficiently small, the acute angle between $\boldsymbol{x}_{2}$ and $V_{A}$ is also sufficiently small. Hence, one can take $\varepsilon$ such that $\boldsymbol{x}_{2}$ and $-\boldsymbol{x}_{1}$ lie in the same cone out of the four cut by $V_{B}$ and $W_{B}$. Furthermore, $\boldsymbol{x}_{2}$ and $\boldsymbol{x}_{1}$ lie in the same cone out of the four cut by $V_{A}$ and $W_{A}$.
@@
    Next consider the case $(\operatorname{Lat}(A)$, $\operatorname{Lat}(B)$, $\operatorname{Lat}(A)$, $\operatorname{Lat}(B))$. Denote by $V_{A},V_{B},W_{A},W_{B}$ these subspaces in this order. Without loss of generality suppose $\boldsymbol{x}\notin V_{B}\cup W_{B}$. The strategy is exactly the same as the previous case (see Figure 8). However, in the present case, we need to apply Lemma 9 for a total of six times, where $V$ will be the subspaces $V_{B},W_{A},W_{B},V_{A},V_{B},W_{A}$ respectively.
@@
    (In fact, the four vectors $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\boldsymbol{x}_{5},\boldsymbol{x}_{6}$ would already suffice). Similarly, since the element $(A,\boldsymbol{a})$ (and hence $(I,\boldsymbol{0})$) can be represented as a full-image word over $\mathcal{G}$, the semigroup $\langle\mathcal{G}\rangle$ is a group by Lemma 1.
@@
Without loss of generality, let $A$ be the shear and $B$ be the positive scale. Starting from any one-dimension subspace $S$, rotate $S$ counter-clockwise and consider its sequence of encounters with $\operatorname{Lat}(A)$ and $\operatorname{Lat}(B)$ before finishing a full cycle. The sequence of encounters must be a cyclic permutation of $(\operatorname{Lat}(A),\operatorname{Lat}(B),\operatorname{Lat}(B))$. Denote by $V_{A},V_{B},W_{B}$ be these subspaces in this order. Without loss of generality suppose $\boldsymbol{x}\notin V_{B}\cup W_{B}$, the case where $\boldsymbol{x}\notin V_{A}$ is analogous.
For a small enough $\varepsilon$, apply Lemma 9 to $\boldsymbol{x}$ to obtain $\boldsymbol{x}_{1}$ that is sufficiently close to $V_{B}$; then apply Lemma 10 to $\boldsymbol{x}_{1}$ to obtain $\boldsymbol{x}_{2}$ that is sufficiently close to $V_{A}$ and therefore lies in the same cone cut by $V_{B}$ and $W_{B}$ as $-\boldsymbol{x}_{1}$; then apply Lemma 9 to $\boldsymbol{x}_{2}$ to obtain $\boldsymbol{x}_{3}$ that is sufficiently close to $V_{B}$ and therefore lies in a different halfspace cut by $V_{A}$ than $\boldsymbol{x}_{2}$; finally, apply Lemma 10 to $\boldsymbol{x}_{3}$ to obtain $\boldsymbol{x}_{4}$ that is sufficiently close to $V_{A}$ and lies in the same cone cut by $V_{B}$ and $W_{B}$ as $-\boldsymbol{x}_{3}$. In this way, we obtain $(I,\boldsymbol{x}_{1}),(I,\boldsymbol{x}_{2}),(I,\boldsymbol{x}_{3}),(I,\boldsymbol{x}_{4})$ with $\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{4}$ generating $\mathbb{Q}^{2}$ as a $\mathbb{Q}_{\geq 0}$-cone. Hence, there exist positive integers $n_{1},\ldots,n_{4}$ such that